# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v4

@@ II-D Some Distributions
  • Gaussian distributions array $P_{l}^{(m)}\sim\mathcal{N}(\bm{0},\mu\Psi(m)\bm{I}_{l})$ on $\mathbb{R}^{l}$ with $10<\mu<$ 1 where $\Psi(m)$ is the variance scaling function. The density function of $P_{l}^{(m)}$ is written as
@@
    $f^{m}_{l}(\bm{x})=\left\{\begin{split}&\frac{1}{\Delta}g^{m}_{l}(\bm{x}),\sqrt{\mu^{2}l\Psi(m)}\leq\|\bm{x}\|\leq\sqrt{l\Psi(m)},\\ &0,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,otherwise\end{split}\right.$ (15)
@@
III Technical Prepartions
For covert communication over AWGN channels, the following distributions are closely related.
For each $n$, the following distributions are defined.
@@ III Technical Prepartions
    $f^{(n)}(\bm{x})=\left\{\begin{split}&\frac{1}{\Delta}g^{(n)}(\bm{x}),\sqrt{\mu^{2}n\Psi(n)}\leq\|\bm{x}\|\leq\sqrt{n\Psi(n)},\\ &0,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,otherwise\end{split}\right.$ (19)
    In (19), we have $\Psi(n)=O(\frac{1}{n^{\tau}})$ with $10<\tau<$ 1, which will be explained later. The variable $\Delta$ is the normalized coefficient and equals to
    $\Delta=E_{P_{n}^{(n)}}[1_{\{\bm{x}\in\mathfrak{B}^{n}_{0}(\sqrt{n\Psi(n)})\backslash\mathfrak{B}^{n}_{0}(\sqrt{\mu^{2}n\Psi(n)})\}}],$ (20)
@@
    The distribution of channel noise at the main channel with blocklength $n$: $Q^{(n)}_{m,0}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$, the probability density function is denoted as $f_{0}^{(n)}(\bm{y})$.
@@
    The output distribution $Q^{(n)}_{m,1}\sim\mathcal{N}(\bm{0},(1+\mu\Psi(n))\bm{I}_{n})$ of the main channel corresponding to the input distribution $P_{n}^{(n)}$. The probability density function of $Q^{(n)}_{m,1}$ is denoted as $g_{1}^{(n)}(\bm{y})$. From the definition of the channel, $g_{1}^{(n)}=g^{(n)}*f_{0}^{(n)}$.
  5. 5.
    The output distribution $\bar{Q}^{(n)}_{m,1}$ of the main channel corresponding to the input distribution $\bar{P}_{n}^{(n)}$. The probability density function of $\bar{Q}^{(n)}_{m,1}$ is denoted as $\bar{g}_{1}^{(n)}(\bm{y})$. From the definition of the channel, we have $\bar{g}^{(n)}_{1}=f^{(n)}*f_{0}^{(n)}$.
  6. 6.
    The distribution of channel noise at the adversary with blocklength $n$: $\hat{Q}^{(n)}_{0}(\bm{z})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I}_{n})$, the probability density function of which is denoted as $\hat{f}_{0}^{(n)}(\bm{z})$.
  7. 7.
    The output distribution $\bar{Q}^{(n)}_{1}$ at the adversary corresponding to the input distribution $P_{n}^{(n)}$. The probability density function of $\bar{Q}^{(n)}_{1}$ is denoted as $\bar{f}_{1}^{(n)}(\bm{z})$. From the definition of the channel, we have $\bar{f}^{(n)}_{1}=g^{(n)}*\hat{f}_{0}^{(n)}$.
  8. 8.
    The output distribution $\hat{Q}^{(n)}_{1}$ of the channel corresponding to the input distribution $\bar{P}_{n}^{(n)}$. The probability density function of $\hat{Q}^{(n)}_{1}$ is denoted as $\hat{f}_{1}^{(n)}(\bm{z})$. From the definition of the channel, we have $\hat{f}_{1}^{(n)}=f^{(n)}*\hat{f}_{0}^{(n)}$.
Note that the distributions $P_{n}^{(n)}$ and $\bar{P}_{n}^{(n)}$ are the special cases of Gaussian distributions array and their truncated version in previous definition.
@@
[…]
The proof can be found in Appendix A. The formula (22) and (24) have been reported in [24].
As $n$ increases, the probability mass of Gaussian distributions $P_{n}^{(n)}$ will increasingly concentrate in the region $\mathfrak{B}^{n}_{0}(\sqrt{n\Psi(n)})\backslash\mathfrak{B}^{n}_{0}(\sqrt{\mu^{2}n\Psi(n)})$, which leads to the sphere hardening effect [30]. With $10<\mu<$ 1 being a constant, $\Delta\rightarrow 1$ as $n\rightarrow\infty$ exponentially and the rapidity depends only on $\mu$. Please see Fig.4 for reference. Thus, we have the following lemma for $\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)})$.
Fig. 2: $\Delta$ tends to $1$ with proper chosen $\mu$.
Lemma 2.
@@
Let $\|\cdot\|_{w}^{*}$ be a metric of weak convergence of probability measures, such as Prokhorov metric [31], we have the following lemma.
@@
The pairs of distributions $\bar{P}_{n}^{(n)}$ and $P_{n}^{(n)}$ satisfies,
$\|P_{n}^{(n)}-\bar{P}_{n}^{(n)}\|_{w}^{*}\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty.$ (26)
@@
From the definition of weak convergence, for any bounded continuous function $f(\bm{x})$ on $\mathbb{R}^{n}$, we have
[…]
From the assumption of $f$ and Lemma 2, we have $\|P_{n}^{(n)}-\bar{P}_{n}^{(n)}\|_{w}^{*}\rightarrow 0$ as $n\rightarrow\infty$. ∎
Different from discrete memoryless channels, the input distribution and the output distributions of AWGN channels are related by the convolution operation with the distribution of channel noise. In the following analysis, we treat the input distributions as convolution operators, and leverage the metric of weak convergence in probability theory to characterize the closeness of the output distributions. The following definition is necessary, which can be found in [32].
@@
The following theorem from [32] (pages 255) is necessary for our discussion.
@@
A sequence of probability distributions $P_{n}$ converges properly (weakly) to a probability distribution $P_{0}$ iff for each $u\in C(\mathbb{R}^{n})$ the convolutions $U_{n}=\mathfrak{F}_{n}u$ converge uniformly to a limit $U=\mathfrak{F}u$.
For the terms $\frac{\bar{Q}^{(n)}_{m,1}(\bm{y})}{Q^{(n)}_{m,1}(\bm{y})}$ and $\frac{\hat{Q}^{(n)}_{1}(\bm{z})}{Q^{(n)}_{1}(\bm{z})}$, we have the following bounds when $n$ is sufficiently large.
@@
For sufficiently large $n$, for any fixed $10<\mu<$ 1 and any $\bm{z}$, we have
$1-\iota\leq\frac{\bar{Q}^{(n)}_{m,1}(\bm{y})}{Q^{(n)}_{m,1}(\bm{y})}=\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\leq 1+\iota,$ (28)
$1-\iota^{\prime}\leq\frac{\hat{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{1}(\bm{z})}=\frac{\hat{f}_{1}^{(n)}(\bm{z})}{\bar{f}^{(n)}_{1}(\bm{z})}\leq 1+\iota^{\prime}.$ (29)
@@
The proof can be found in Appendix B. The Lemma is an application of Theorem 1 with the specific distributions $P_{n}^{(n)}$ and $\bar{P}_{n}^{(n)}$.
@@
Note that there is a bit abuse of notations. We don’t distinguish $\bar{Q}^{(n)}_{m,1}(\bm{y})$, $Q^{(n)}_{m,1}(\bm{y})$, $\hat{Q}^{(n)}_{1}(\bm{z})$ and $\bar{Q}^{(n)}_{1}(\bm{z})$ from their probability density functions in (28) and (29). These expressions make our results consistent with that of DMCs and avoid utilization of too many symbols. In the following sections, we will continue to adopt these expressions for convenience.
IV Covert Communication over AWGN Channels
This section provides main results - the first and second order asymptotics of the maximal throughput of covert communication over AWGN channels. The results are organized as follows. In Section IV-A, we present the random coding scheme, the decoding method and the analysis of decoding error probability. In Section IV-B, channel resolvability results are developed.
IV-A Reliability Analysis
This section present one-shot achievability results. The proofs follow the same line as [23]. Nevertheless, as the generating distribution are truncated Gaussian, whose support set is a subset of $\mathbb{R}^{n}$, the error probability of decoding should be considered separately for codewords with different radius in $\mathbb{R}^{n}$. However, by proper choosing the decoding threshold, we ensure that the error probability is small.
The transmitter Alice generates a family of $K$ codebooks $\{\bm{x}_{sw},s\in\langle K\rangle,w\in\langle M\rangle\}$ from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$ by independent and identically distributed sampling. Then, she choose an index $s$ from $\langle K\rangle$. The unique key $\mathsf{S}(s)$ is sent to the legitimate receiver Bob and is unknown to the adversary. With the key $\mathsf{S}(s)$ , Alice choose the $s$-th codebook to communicate with Bob. Meanwhile, Bob utilizes the $s$-th codebook in the threshold decoding upon observing a channel output vector $\bm{y}$. With a parameter $\gamma$ and a fixed distribution $Q^{(n)}_{m,1}$ on $\mathcal{R}^{n}$, the decoder decides $\widehat{W}=w$ as the transmitted message if $w$ is the unique index in $\langle M\rangle$ such that
$\log\frac{W_{1}^{n}(\bm{y}|{\bm{x}_{sw}})}{Q^{(n)}_{m,1}(\bm{y})}>\gamma.$ (30)
If there is no $w$ satisfying (30), the decoder decares an error. This decoder has been adopted in the one shot achievability analysis of [23] and is based on the threshold decoder [2]. Since the candidates of the decoder are within the $s$-th codebook, we write $\bm{x}_{w}$ to replace $\bm{x}_{sw}$ for convenience.
@@ IV-A Reliability Analysis
$\varepsilon_{w}^{(1)}=\mathbb{P}\left(\log\frac{W^{n}(\bm{Y}|{\bm{x}})}{Q^{(n)}_{m,1}(\bm{Y})}\leq\gamma\arrowvert\bm{x}=\bm{x}_{w}\right),$ (31)
$\varepsilon_{w}^{(2)}=\mathbb{P}\left(\exists w^{\prime}\neq w:\log\frac{W^{n}(\bm{Y}|{\bm{x}_{w^{\prime}}})}{Q^{(n)}_{m,1}(\bm{Y})}>\gamma\right).$ (32)
Note that $Q^{(n)}_{m,1}$ is the induced distribution of the channel corresponding to $P_{n}^{(n)}$, which is different from the induced distribution $\bar{Q}^{(n)}_{m,1}$ of the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$. The following lemma is almost a restatement of Lemma 1 in [23]. The differences are: (1) The discrete memoryless channel is substituted by the AWGN channel. (2) The sum is substituted by integration. (3) We use $\log$ with base $2$ instead of base $e$.
Lemma 4.
$\mathbb{E}\left(\varepsilon_{w}^{(1)}\right)=P_{X^{n}}W(Y^{n}|X^{n})\left(\log\frac{W(Y^{n}|X^{n})}{Q_{Y^{n}}}\leq\gamma\right),$ (33)
$\mathbb{E}\left(\varepsilon_{w}^{(2)}\right)=M\cdot 2^{-\gamma}\mathbb{E}_{P_{Y^{n}}}\big{(}\frac{P_{Y^{n}}}{Q_{Y^{n}}}\big{)}.$ (34)
Applying Lemma 4 in our setting, we have
$\mathbb{E}\left(\varepsilon_{w}^{(1)}\right)=\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})W_{1}^{n}(\bm{Y}|\bm{X})}\left(\log\frac{W_{1}^{n}(\bm{Y}|{\bm{X}})}{{g}_{1}^{(n)}(\bm{Y})}\leq\gamma\right),$ (35)
$\mathbb{E}\left(\epsilon_{w}^{(2)}\right)=M_{n}\cdot 2^{-\gamma}\mathbb{E}_{\bar{Q}^{(n)}_{m,1}}\big{(}\frac{\bar{Q}^{(n)}_{m,1}}{Q^{(n)}_{m,1}}\big{)}.$ (36)
Let’s consider the error probability of the first type in (35) firstly, and we have the following lemma.
@@
$\gamma=$ $nC(\mu,\Psi(n))+\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))$
$+$ $\sqrt{n\mathsf{V}(\mu^{2}\Psi(n),\mu\Psi(n))}$
$\times$ $\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})$ (37)
@@
and
$\mathsf{V}(\mu^{2}\Psi(n),\mu\Psi(n))=\frac{2\mu^{2}\Psi(n)+\mu^{2}\Psi^{2}(n)}{2(1+\mu\Psi^{2}(n)}\log^{2}e,$ (40)
then we have
@@
The average decoding error over the generating distribution $\bar{P}_{n}^{(n)}$ in (35) can be rewritten as
$\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})}\left[\mathbb{P}_{W_{1}^{n}(\bm{Y}|\bm{X}=\bm{x})}\bigg{(}\log\frac{W_{1}^{n}(\bm{Y}|{\bm{X}=\bm{x}})}{{g}_{1}^{(n)}(\bm{Y})}\leq\gamma\bigg{)}\right].$ (42)
For a codeword $\bm{x}$ sampled from truncation Gaussian distribution $\bar{P}_{n}^{(n)}$, we consider the probability of error in (31). Due to the sphere symmetry, we assume $\bm{x}_{0}=[\sqrt{\Upsilon},\cdots,\sqrt{\Upsilon}]$, where $\Upsilon\in[\mu^{2}\Psi,\Psi]$. The information density $\log\frac{W_{1}^{n}(\bm{y}|{\bm{x}_{0}})}{{g}_{1}^{(n)}(\bm{y})}$ is expressed as
$\frac{n}{2}\log(1+\mu\Psi(n))+\frac{\log e}{2}\sum_{i=1}^{n}\left[\frac{y_{i}^{2}}{1+\mu\Psi(n)}-(y_{i}-\sqrt{\Upsilon})^{2}\right].$ (43)
@@
$\begin{split}H^{\Upsilon}_{n}=&\frac{n}{2}\log(1+\mu\Psi(n))\\ +&\frac{\log e}{2(1+\mu\Psi(n))}\sum_{i=1}^{n}\left[\Upsilon-\mu\Psi(n)\xi_{i}^{2}+2\sqrt{\Upsilon}\xi_{i}\right]\end{split}$ (44)
@@
$\begin{split}\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n))\right]=&\frac{\Upsilon-\mu\Psi(n)}{2(1+\mu\Psi(n))}\log e\\ =&\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n)),\end{split}$ (46)
$\begin{split}\mathsf{Var}\left[\frac{1}{n}\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n))\right]=&\frac{1}{n}\frac{2\Upsilon+\mu^{2}\Psi^{2}}{2(1+\mu\Psi)^{2}}\log^{2}e\\ =&\frac{\mathsf{V}(\Upsilon,\mu\Psi(n))}{n},\end{split}$ (47)

### v5

@@ II-E Related Distributions
  • Gaussian distributions array $P_{l}^{(m)}\sim\mathcal{N}(\bm{0},\mu\Psi(m)\bm{I}_{l})$ on $\mathbb{R}^{l}$ with $0<\mu<1$ where $\Psi(m)$ is the variance scaling function. The density function of $P_{l}^{(m)}$ is written as
@@
    $\bar{g}^{m}_{l}(\bm{x})=\left\{\begin{split}&\frac{1}{\Delta}g^{m}_{l}(\bm{x}),\sqrt{\mu^{2}l\Psi(m)}\leq\hskip-2.8903pt\|\bm{x}\|\hskip-2.8903pt\leq\hskip-2.8903pt\sqrt{l\Psi(m)},\\ &0,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,otherwise\end{split}\right$. (21)
@@
For covert communication over AWGN channels, the following distributions are closely related. We introduce them according to where they lie at.
  • Distributions at the transmitter
@@
      $\bar{g}^{(n)}(\bm{x})=\left\{\begin{split}&\frac{1}{\Delta}g^{(n)}(\bm{x}),\sqrt{\mu^{2}n\Psi(n)}\leq\hskip-2.8903pt\|\bm{x}\|\hskip-2.8903pt\leq\hskip-2.8903pt\sqrt{n\Psi(n)},\\ &0,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,otherwise\end{split}\right$. (25)
      with $0<\mu<1$. The variable $\Delta$ is the normalized coefficient and equals to
      $\Delta=E_{P_{n}^{(n)}}[1_{\{\bm{x}\in\mathfrak{B}^{n}_{0}\big(\sqrt{n\Psi(n)}\big)\backslash\mathfrak{B}^{n}_{0}\big(\sqrt{\mu^{2}n\Psi(n)}\big)\}}],$ (26)
@@
  • Distributions at the legitimate receiver Bob.
    1. 1.
      The distribution of channel noise at the main channel with blocklength $n$: $Q^{(n)}_{\bm{Y},0}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$, the probability density function is denoted as $f_{0}^{(n)}(\bm{y})$.
    2. 2.
      The output distribution $Q^{(n)}_{\bm{Y},1}\sim\mathcal{N}(\bm{0},(1+\mu\Psi(n))\bm{I}_{n})$ of the main channel corresponding to the input distribution $P_{n}^{(n)}$. The probability density function of $Q^{(n)}_{\bm{Y},1}$ is denoted as $f_{1}^{(n)}(\bm{y})$. From the definition of the channel, $f_{1}^{(n)}=g^{(n)}*f_{0}^{(n)}$.
@@
      The output distribution $\bar{Q}^{(n)}_{\bm{Y},1}$ of the main channel corresponding to the input distribution $\bar{P}_{n}^{(n)}$. The probability density function of $\bar{Q}^{(n)}_{\bm{Y},1}$ is denoted as $\bar{g}_{1}^{(n)}(\bm{y})$. From the definition of the channel, we have $\bar{f}^{(n)}_{1}=\bar{g}^{(n)}*f_{0}^{(n)}$.
  • Distributions at the adversary Willie.
    1. 1.
      The distribution of channel noise at the adversary with blocklength $n$: $Q^{(n)}_{\bm{Z},0}(\bm{z})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I}_{n})$, the probability density function of which is denoted as $h_{0}^{(n)}(\bm{z})$.
    2. 2.
      The mixture Gaussian distribution induced by the codebook at the adversary $Q^{(n)}_{\bm{Z},\bm{C}}$, which is defined in (13).
    3. 3.
      The output distribution $Q^{(n)}_{\bm{Z},1}$ at the adversary corresponding to the input distribution $P_{n}^{(n)}$. The probability density function of $Q^{(n)}_{\bm{Z},1}$ is denoted as $h_{1}^{(n)}(\bm{z})$. From the definition of the channel, we have $h^{(n)}_{1}=g^{(n)}*h_{0}^{(n)}$.
@@
      The output distribution $\bar{Q}^{(n)}_{\bm{Z},1}$ of the channel corresponding to the input distribution $\bar{P}_{n}^{(n)}$. The probability density function of $\bar{Q}^{(n)}_{\bm{Z},1}$ is denoted as $\bar{h}_{1}^{(n)}(\bm{z})$. From the definition of the channel, we have $\bar{h}_{1}^{(n)}=\bar{g}^{(n)}*h_{0}^{(n)}$.
Note that the distributions $P_{n}^{(n)}$ and $\bar{P}_{n}^{(n)}$ are the special cases of Gaussian distributions array and their truncated version in previous definition. The next lemma and two corollaries characterize the closeness of the distributions $P_{n}^{(n)}$ and $\bar{P}_{n}^{(n)}$.
@@
$\gamma(a,z)=\int_{0}^{z}e^{-t}t^{a-1}dt.$ (30)
The proof can be found in Appendix B. With $0<\mu<1$ being a constant, $\Delta\rightarrow 1$ as $n\rightarrow\infty$ exponentially and the rapidity depends only on $\mu$. Please see Fig.2 for reference. Thus, we have the following corollary for $\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)})$, the details of the analysis has been presented in Section IV-A of [23].
Fig. 2: $\Delta$ tends to 11 with proper chosen $\mu$.
Corollary 1.
@@
Let $C_{0}(-\infty,\infty)$ be the class of bounded continuous functions vanishing at $\infty$ and $C[-\infty,\infty]$ be the class of continuous functions whose limits at -$\infty$ and $\infty$ are finite, which are both the subclasses of bounded continuous functions and have relationship as $C_{0}(-\infty,\infty)\subset C[-\infty,\infty]$. We first recall the following theorem (Theorem 1 at page 249 of [36]).
Theorem 1.
In order that a probability distribution $F_{n}$ converges properly [^1] Proper convergence is equivalent to weak convergence with respect to $C_{0}(-\infty,\infty)$. to $F$, it is necessary and sufficient that
$\mathbb{E}_{n}(u)\rightarrow\mathbb{E}(u)\,\ \ \,\text{for all}\,\ \ u\in C_{0}(-\infty,\infty)$ (32)
where $\mathbb{E}_{n}$ and $\mathbb{E}$ are expectation with respect to $F_{n}$ and $F$.
Let $\|\cdot\|_{m}^{*}$ be a metric of proper convergence of probability measures, we have the following corollary.
@@
The distribution pair $\bar{P}_{n}^{(n)}$ and $P_{n}^{(n)}$ satisfies
$\|\bar{P}_{n}^{(n)}-P_{n}^{(n)}\|_{m}^{*}\rightarrow 0.$ (33)
@@
For any continuous function $u(\bm{x})$ on $\mathbb{R}^{n}$ which vanishes at infinity, we have
$\begin{split}&|\int u(\bm{x})P_{n}^{n}(d\bm{x})-\int u(\bm{x})\bar{P}_{n}^{n}(d\bm{x})|\\ \leq&\int u(\bm{x})|g^{n}(\bm{x})-\bar{g}^{n}(\bm{x})|d\bm{x}\\ \leq&\,\,2\,\,\underset{\bm{x}}{\sup}|u(\bm{x})|\cdot\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)}).\end{split}$
From the assumption of $u$ and Corollary 1, we have
$\mathbb{E}_{\bar{P}_{n}^{(n)}}(u)\rightarrow\mathbb{E}_{P_{n}^{(n)}}(u).$ (34)
By using a multivariate version of Theorem 1, we have the conclusion. ∎
Different from discrete memoryless channels, the input distribution and the output distributions of AWGN channels are related by the convolution operation with the distribution of channel noise. In the following analysis, we treat the input distributions as convolution operators, and leverage the metric of proper convergence in probability theory to characterize the closeness of the output distributions at both the main channel $W_{1}$ and the channel at the adversary $W_{2}$. The following definition is necessary, which can be found in [36].
@@
The following theorem from [36] (page 255) is necessary for our discussion.
@@
A sequence of probability distributions $P_{n}$ converges properly to a probability distribution $P_{0}$ iff for each $u\in C[-\infty,\infty]$, the convolutions $U_{n}=\mathfrak{F}_{n}u$ converge uniformly to a limit $U=\mathfrak{F}u$.
For the terms $\frac{\bar{Q}^{(n)}_{\bm{Y},1}(\bm{y})}{Q^{(n)}_{\bm{Y},1}(\bm{y})}$ and $\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}$, we have the following bounds when $n$ is sufficiently large.
@@
For sufficiently large $n$, for any fixed $0<\mu<1$ and any $\bm{z}$, we have
$1-\iota\leq\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\leq 1+\iota,$ (36)
$1-\iota^{\prime}\leq\frac{\bar{h}_{1}^{(n)}(\bm{z})}{h^{(n)}_{1}(\bm{z})}\leq 1+\iota^{\prime}.$ (37)
@@
The proof can be found in Appendix C and is an application of Theorem 2 with the specific distributions $P_{n}^{(n)}$ and $\bar{P}_{n}^{(n)}$.
@@
In the follow-up analysis, we don’t distinguish $\bar{Q}^{(n)}_{\bm{Y},1}(\bm{y})$, $Q^{(n)}_{\bm{Y},1}(\bm{y})$, $\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})$ and $Q^{(n)}_{\bm{Z},1}(\bm{z})$ from their probability density functions. These expressions make our results consistent with that of DMCs and avoid utilization of too many symbols. In the following sections, we will continue to adopt these expressions for convenience.
III Throughput Analysis of Covert Communications Over AWGN channels
This section provides main results - the first and second order asymptotics of the maximal throughput of covert communication over AWGN channels. The results are organized as follows. In Section III-A, we present the random coding scheme, the decoding method and the analysis of decoding error probability. In Section III-B, channel resolvability results are developed.
III-A Random Coding, Decoding and Reliability Analysis
This section present one-shot achievability results. This part differs from previous works over DMCs on the choices of the generating distribution and the reference distribution for decoding. The proofs follow the same line as [22] by considering the two types of error $\varepsilon_{sw}^{(1)}$ and $\varepsilon_{sw}^{(2)}$ separately. Nevertheless, as the generating distribution are truncated Gaussian, whose support set is a subset of $\mathbb{R}^{n}$, the error probability $\varepsilon_{sw}^{(1)}$ can be considered separately for codewords with different radius in $\mathbb{R}^{n}$. By properly choosing the decoding threshold, we ensure that $\varepsilon_{sw}^{(1)}$ for each $(s,w)$ pair is small. For error in the second type $\varepsilon_{sw}^{(2)}$, we bound the expectation of it for the generation distribution, and we will use expurgation operation later to pick part of codewords such that this type of errors of them are also small.
The transmitter Alice generates a family of $K$ codebooks $\{\bm{x}_{sw},s\in\langle K\rangle,w\in\langle M\rangle\}$ from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$ by independent and identically distributed sampling. Then, she choose an index $s$ from $\langle K\rangle$. The unique key $\mathsf{S}(s)$ is sent to the legitimate receiver Bob and is unknown to the adversary. With the key $\mathsf{S}(s)$ , Alice choose the $s$-th codebook to communicate with Bob. Meanwhile, Bob utilizes the $s$-th codebook in the threshold decoding upon observing a channel output vector $\bm{y}$. With a parameter $\gamma$ and a fixed distribution $Q^{(n)}_{\bm{Y},1}$ on $\mathcal{R}^{n}$, the decoder decides $\widehat{W}=w$ as the transmitted message if $w$ is the unique index in $\langle M\rangle$ such that
$\log\frac{W_{1}^{n}(\bm{y}|{\bm{x}_{sw}})}{Q^{(n)}_{\bm{Y},1}(\bm{y})}>\gamma.$ (38)
If there is no $w$ satisfying (38), the decoder decares an error. This decoder has been adopted in the one shot achievability analysis of [22] and is based on the threshold decoder [20].
@@ III-A Random Coding, Decoding and Reliability Analysis
[…]
$\varepsilon_{sw}^{(2)}=\mathbb{P}\left(\exists{w}^{\prime}\neq w:\log\frac{W_{1}^{n}(\bm{Y}|{\bm{x}_{sw^{\prime}}})}{Q^{(n)}_{\bm{Y},1}(\bm{Y})}>\gamma\right).$ (40)
Note that $Q^{(n)}_{\bm{Y},1}$ is the induced distribution of the channel corresponding to $P_{n}^{(n)}$, which is different from the induced distribution $\bar{Q}^{(n)}_{\bm{Y},1}$ of the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$.
As maximal error probability is the reliability metric, we consider the two types of error $\varepsilon_{sw}^{(1)}$ and $\varepsilon_{sw}^{(2)}$ separately. The bounding of the first type of error will make use of the Berry-Esseen Theorem, which we recall here for convenience.
Theorem 3.
(Berry-Esseen Theorem): Let $X_{1},\cdots,X_{n}$ be independent random variables such that for $i\in\langle n\rangle$, $\mathbb{E}(X_{i})=u_{i}$, $Var(X_{i})=v_{i}^{2}$ and $t_{i}=\mathbb{E}\big(|X_{i}-\mu_{i}|^{3}\big)$. If we define $v^{2}=\sum_{i=1}^{n}v_{i}^{2}$ and $t=\sum_{i=1}^{n}t_{i}$, then we have
$|\mathbb{P}\bigg(\sum_{i=1}^{n}(X_{i}-u_{i})\leq\lambda v\bigg)-\Phi(\lambda)|\leq\frac{6t}{v^{3}}.$ (41)
The following lemma bounds the maximum of $\varepsilon_{sw}^{(1)}$ by choosing proper decoding threshold $\gamma$. As the power of the codewords ranges from $\mu^{2}n\Psi(n)$ to $n\Psi(n)$, the bounding of the error is slightly different from previous works [20][24], where sphere symmetry can be leveraged for codewords with equal power. Nevertheless, by separately considering the codewords with equal power and choosing an universal decoding threshold, the analysis can also be simplified by leveraging sphere symmetry.
@@
$\gamma=$ $nC(\mu,\Psi(n))+n\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))$
$+$ $\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}$
$\times$ $\Phi^{-1}\bigg(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\bigg)$ (42)
@@
$\mathsf{V}(\Psi(n),\mu\Psi(n))=\frac{2\Psi(n)+\mu^{2}\Psi^{2}(n)}{2(1+\mu\Psi^{2}(n))}\log^{2}e,$ (45)
and $\Phi$ is the cumulative distribution function of standard normal random variable, then we have
@@
For a codeword $\bm{x}_{sw}$ being sampled from truncation Gaussian distribution $\bar{P}_{n}^{(n)}$, let $\|\bm{x}_{sw}\|^{2}=n\Upsilon$, where $\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$. We consider the probability of error in (39) for $\bm{x}_{sw}$ on the n1n-1 dimensional sphere with radius $\sqrt{n\Upsilon}$. Due to the sphere symmetry of the reference distribution $Q^{(n)}_{\bm{Y},1}$, we assume $\bm{x}_{0}=[\sqrt{\Upsilon},\cdots,\sqrt{\Upsilon}]$. The sphere symmetry has been employed in previous works [24] and [20] to simplify the analysis of decoding error. The information density $\log\frac{W_{1}^{n}(\bm{Y}|{\bm{x}_{0}})}{Q^{(n)}_{\bm{Y},1}(\bm{Y})}=\log\frac{W_{1}^{n}(\bm{Y}|{\bm{x}_{0}})}{f_{1}^{(n)}(\bm{y})}$ is expressed as
$\frac{n}{2}\log\big(1+\mu\Psi(n)\big)+\frac{\log e}{2}\sum_{i=1}^{n}\left[\frac{y_{i}^{2}}{1+\mu\Psi(n)}-(y_{i}-\sqrt{\Upsilon})^{2}\right].$ (47)
@@
$\begin{split}H^{\Upsilon}_{n}=&\frac{n}{2}\log\big(1+\mu\Psi(n)\big)\\ +&\frac{\log e}{2(1+\mu\Psi(n))}\sum_{i=1}^{n}\left[\Upsilon-\mu\Psi(n)\xi_{i}^{2}+2\sqrt{\Upsilon}\xi_{i}\right]\end{split}$ (48)
@@
$\begin{split}\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n))\right]=&\frac{\Upsilon-\mu\Psi(n)}{2(1+\mu\Psi(n))}\log e\\ =&\bar{\mathsf{C}}\big(\Upsilon,\mu\Psi(n)\big),\end{split}$ (50)
$\begin{split}\mathsf{Var}\left[\frac{1}{n}\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n))\right]=&\frac{1}{n}\frac{2\Upsilon+\mu^{2}\Psi^{2}}{2(1+\mu\Psi)^{2}}\log^{2}e\\ =&\frac{\mathsf{V}\big(\Upsilon,\mu\Psi(n)\big)}{n},\end{split}$ (51)